I'm having trouble solving circumference ratios. The question is....
A boy rides his bicycle around a circular track that has a diameter of 76 ft. If a wheel of the bicycle spins exactly 48 times when going once around the track, what is the diameter of the wheel?
Write your answer in inches.
The answer key gives this formula to follow: n=D3.14/d3.14 =D/d
And here are the steps they made, but I don't know why they multiplied 76 by 12.
n=48
D=76=76*12=912
d=912/48=19
d=19inches

Found 2 solutions by Abbey, cleomenius:
Answer by Abbey(339)

You have to have the same measurements when solving word problems.
in order to convert 76 feet into inches, we multiply by 12 (12 inches per foot)

Answer by cleomenius(959)

First step I did was find the convert to inches the diameter of the track.
76 * 12 = 912
Next, find the length of the track.
3.14 * 912 = 2865
Next, divide the distance by the number of revolutions of the wheel.
59.7
This is the distance traveled by one revolution of the wheel.
This distance = pi * the diameter of the wheel.
59.7 / pi = 19 inches for the diameter of the wheel.
The reason they multiplied by 12 is to have the calculations of the equation put into inches instead of feet.